# A high-order procedure for computing globally optimal Wannier functions in one-dimensional crystalline insulators

### v2

@@
A standard task in solid state physics and quantum chemistry is the computation of localized molecular orbits known as Wannier functions. In this manuscript, we propose a new procedure for computing Wannier functions in one-dimensional crystalline materials. Our approach proceeds by first performing parallel transport of the Bloch functions using numerical integration. Then, using novel analysis, we show that a simple correction can be analytically computed that yields the optimally localized Wannier function. The resulting scheme is robust and capable of achieving high-order accuracy. We illustrate this in a number of numerical experiments.
A high-order procedure for computing globally optimal Wannier functions in one-dimensional crystalline insulators
@@ 1 Introduction
We now outline the remainder of this manuscript. In Section 2, we review some background material on Wannier functions. This is followed by Section 3 where we develop the analytic apparatus for this work. After this are Sections 4.1 and 5 where we describe our numerical procedure. Our numerical procedure is illustrated in Section 6 which contains numerical experiments. We provide our conclusions and outline future work in Section 7.
@@ 2.1 Notation
We let $\ell^{2}$ denote the Hilbert space for complex sequences $\mathbf{x}=\left\{x_{i}\right\}_{i=-\infty}^{\infty}$ such that $\sum_{i=-\infty}^{\infty}\absolutevalue{x_{i}}^{2}<\infty$ with the standard inner product. We denote the right shift operator in $\ell^{2}$ by $R$, where
@@
For a vector $\mathbf{x}\in\ell^{2}$ or $\mathbb{C}^{N}$, we denote by $\mathbf{x}^{*}$ the conjugate transpose and by $\|\mathbf{x}\|$ the standard 2-norm, $\|\mathbf{x}\|=\sqrt{\mathbf{x}^{*}\mathbf{x}}$.
Given sets $A\subset\mathbb{C}$ and $B\subset\mathbb{C}$, we write $f\in C(A;B)$ if $f:A\to B$ is a continuous function. Similarly, we write $f\in C^{n}(A;B)$ if $f$ is a $n$-times continuously differentiable. For an $\mathbf{A}\in\mathbb{C}^{M\times M}$ matrix, we let $\mathbf{A}^{\dagger}$ denote its Moore–Penrose pseudoinverse.
@@ 2.2 Bloch functions
$\psi_{k}^{(j)}(x+y)=e^{iky}\psi_{k}^{(j)}(x),\quad y\in\Lambda.$ (10)
The value $k$ is called the quasimomentum, $\{\psi_{k}^{(j)}\}_{j=1}^{\infty}$ in (9) are called the Bloch functions, and $\{E_{k}^{(j)}\}_{j=1}^{\infty}$ are called the energy states.
@@
$\left[-\left(\frac{\mathrm{d}}{\mathrm{d}x}+ik\right)^{2}+V\right]u_{k}^{(j)}=E_{k}^{(j)}u_{k}^{(j)},$ (13)
@@
$H(k)=-\left(\frac{\mathrm{d}}{\mathrm{d}x}+ik\right)^{2}+V.$ (14)
@@ 2.3 Wannier functions
$W_{0}^{(j)}(x)=\frac{1}{\Omega}\int_{-\Omega/2}^{\Omega/2}\psi_{k}^{(j)}(x)\,{\rm d}k=\frac{1}{\Omega}\int_{-\Omega/2}^{\Omega/2}e^{ixk}u_{k}^{(j)}(x)\,{\rm d}k,\quad j=1,2,\ldots.$ (22)
It was shown in [14], in the case of a symmetric potential, and in [9], more generally, that $\{\psi_{k}^{(j)}\}_{j=1}^{\infty}$ can be scaled such that the $\{W_{0}^{(j)}\}_{j=1}^{\infty}$ are exponentially localized. In other words, there exist constants $C>0$ and $D>0$ such that
@@ 2.4 Analyticity of eigenvalues and eigenvectors
Theorem 2.3 concerns only the eigenprojectors, not the eigenvectors; this is because for any an eigenvector $\mathbf{v}_{i}(z)$ of $T(z)$, one can multiply the vector by a $z$-dependent phase $e^{-\mathrm{i}\phi(z)}$ ($\phi(z)$ is real for real $z$) while the new vector $e^{-\mathrm{i}\phi(z)}\mathbf{v}_{i}(z)$ is still an eigenvector. The phase could change the smoothness of $\mathbf{v}_{i}(z)$ while the projector $P_{i}(z)=\mathbf{v}_{i}(z)\mathbf{v}^{*}_{i}(z)$ is unchanged for real $z$.
@@ 2.5 Derivatives of eigenvalues and eigenvectors
Next, we provide expressions for the derivatives of eigenvalues and eigenvectors. For simplicity, we only state the result for the Hilbert space $\ell^{2}$ and operators with a purely discrete spectrum, which will suffice for our purpose. These formulas can be found in Section 8.2.3 in [13]. The formula (28) is a consequence of the derivative of the projection $\frac{\mathrm{d}P}{\mathrm{d}z}$ substituted into (26).
@@
Suppose that $T(z)$ is a self-adjoint analytic family of operators in $\ell^{2}$ for $z\in D$ and the operators have a purely discrete spectrum. Suppose further that $z_{0}\in D$ is real and $\lambda(z_{0})$ is a nondegenerate eigenvalue of $T(z_{0})$ with the eigenvector $\mathbf{v}$, i.e. $T(z_{0})\mathbf{v}(z_{0})=\lambda(z_{0})\mathbf{v}(z_{0})$. The derivative of $\lambda$ at $z_{0}$ is given by
@@ 3.1 Analysis of Bloch functions
$u(x,k+g)=e^{-igx}u(x,k),\quad g\in\Lambda^{*},\,x\in I_{\mathrm{uc}}.$ (34)
@@
The following lemma provides a simple relation for the dependence of $k$ and $m$ in the Fourier coefficients in (35). It is a well-known relation in the context of periodic zone scheme (see (17)-(19))). It is a simple consequence of (34).
@@
$a_{m}(k)=\alpha(k+m\Omega),\quad k\in\mathbb{R},\,m\in\mathbb{Z}.$ (39)
@@
Lemma 39 shows that we can obtain the function $\alpha$ defined in $\mathbb{R}$ by “unfolding” the Fourier coefficients $a_{m}(k)$ defined in $I_{\mathrm{bz}}$ for $m\in\mathbb{Z}$ such that $a_{m}(k)$ takes up the interval $[m-\Omega/2,m+\Omega/2)$. We summarize this fact in the following lemma, together with Lemma 3.3 and the decay condition (38) in terms of $\alpha$.
@@
where $\alpha$ is defined in (39). The function $\alpha(k)$ can be chosen to be analytic in the intervals $[m-\Omega/2,m+\Omega/2)$ for $m\in\mathbb{Z}$. Furthermore, |$\absolutevalue{\alpha(k)}$ | decays no slower than $|k|^{-2}$ as $|k|\to\infty$.
@@
Since $u(x,k)$ may not be periodically analytic in $k\in I_{\mathrm{bz}}$ (see Remark 3.2), the function $\alpha$ can be a discontinuous function in $\mathbb{R}$. Lemma 3.3 only ensures $\alpha$ is analytic in each interval $[m-\Omega/2,m+\Omega/2)$ for $m\in\mathbb{Z}$. Moreover, the function $\alpha$ is formed from a very special choice of coefficients $a_{m}$ since the corresponding $u(x,k)$ is chosen to be analytic in $k$; this does not happen if one computes $u(x,k)$ at each point $k$ independently (see Remark 2.5) so that $u(x,k)$ is an arbitrarily irregular function in $k$.
@@ 3.2 Properties of Wannier functions
$W_{0}(x)=\frac{1}{\Omega}\int_{-\Omega/2}^{\Omega/2}\sum_{m=-\infty}^{\infty}\alpha(k+m\Omega)e^{i(k+m\Omega)x}\mathrm{d}k=\frac{1}{2\pi}\int_{-\infty}^{\infty}\frac{2\pi}{\Omega}\alpha(\xi)e^{\mathrm{i}\xi x}\mathrm{d}\xi.$ (47)
@@
$\widehat{W}_{0}(\xi)=\int_{-\infty}^{\infty}W_{0}(x)e^{-i\xi x}\,{\rm d}x=\frac{2\pi}{\Omega}\alpha(\xi).$ (49)
@@ 3.3 Perturbation analysis in the Fourier domain
where the vector with elements $\{\widehat{V}_{l}\}_{l=-\infty}^{\infty}$ is in $\ell^{2}$ for piecewise continuous $V$ in $I_{\mathrm{uc}}$. Moreover, we define a vector $\mathbf{y}$ containing the Fourier coefficients of $u$ in (35) and (45):
@@
It follows that $\mathbf{y}\in\ell^{2}$ for any $u$ that satisfies the assumptions of Lemma 3.5. Moreover, it follows from (37) that
@@
We observe that the operator $-\mathbf{\Theta}^{\dagger}(k)\mathbf{S}(k)$ in (61) has a spectrum that decays as $1/m$ asymptotically for any $k\in I_{\mathrm{bz}}$, so it is a compact operator on $\ell^{2}$. As a result, despite the fact that (61) is infinite-dimensional, (61) is very benign both mathematically and numerically.
@@ 3.4 Analysis of the discontinuities at zone boundaries
As discussed in Remark 3.6, $a_{m}(\Omega/2)$ is not guaranteed to be equal to $a_{m+1}(-\Omega/2)$. Since $a_{m}(\Omega/2)$ and $a_{m+1}(-\Omega/2)$ may be viewed as limits of $\alpha(k)$ at the boundary points $k=m-\Omega/2$ for $m\in\mathbb{Z}$ from both sides, their differences characterize the jump of $\alpha(k)$ at these boundary points. The following lemma shows that $a_{m}(\Omega/2)$ and $a_{m+1}(-\Omega/2)$ only differ by a phase independent of $m$.
@@ 3.7 Realty of Wannier functions
While we have completed the construction of optimally localized Wannier functions, there is one remaining degree of freedom, the phase choice of the vector $\mathbf{y}_{0}$ in (63); we could replace $\mathbf{y}_{0}$ by $e^{-\mathrm{i}\phi_{0}}\mathbf{y}_{0}$ for any real $\phi_{0}$ and all above formulas would still hold, including the expressions in Corollary 3.14. However, the phase $e^{-\mathrm{i}\phi_{0}}$ does change the resulting $W_{0}$. In this section, we prove that the constructed Wannier function $W_{0}$ can always be chosen to be real if $\mathbf{y}_{0}$ in (63) satisfies a simple condition.
@@
We observe that the condition in (100) is applicable after solving (61) for determining $\mathbf{y}$ in $[-\Omega/2,\Omega/2]$; we can obtain a constant phase $e^{-\mathrm{i}\phi_{0}}$ to enforce (100) so the the resulting $W_{0}$ is real.
@@ 6 Numerical experiments
We implemented the procedure described in Section 4.1 in Fortran. We fix $V$, $a$, $M$, and $\ell$ and increase $K$ until 10 digits of accuracy is obtained. The timings reported are the CPU times required to obtain the Fourier transform of $W_{0}$ evaluated on a uniform grid (i.e., Steps 1–5 in Section 5). To obtain an estimate of the error we compute
@@ 7 Conclusions
Since Kato’s theorem (Theorem 2.3) guarantees analyticity even when eigenvalues become degenerate in one dimension, the approach in this paper has been easily extended to isolated multiband cases in one dimension. Similarly, globally optimal solutions can be computed analytically. For higher dimensions, topological phenomena show up in the construction of Wannier functions and new ideas beyond this paper are required. Nonetheless, the algorithms in one dimension are the basic building block for higher-dimensional cases, and the extension of the approach in this paper has been worked out and will be reported at a later date.

### v3

# A highly accurate procedure for computing globally optimal Wannier functions in one-dimensional crystalline insulators

@@
A standard task in solid state physics and quantum chemistry is the computation of localized molecular orbitals known as Wannier functions. In this manuscript, we propose a new procedure for computing Wannier functions in one-dimensional crystalline materials. Our approach proceeds by first performing parallel transport of the Bloch functions using numerical integration. Then a simple analytically computable correction is introduced to yield the optimally localized Wannier function. The resulting scheme is rapidly convergent and proven to produce globally optimal Wannier functions. The analysis in this manuscript can also be viewed as a proof of the existence of exponentially localized Wannier functions in one dimension. We illustrate the performance of the scheme by a number of numerical experiments.
A highly accurate procedure for computing globally optimal Wannier functions in one-dimensional crystalline insulators
@@ 1 Introduction
We now outline the remainder of this manuscript. In Section 2, we review some background material on Wannier functions. This is followed by Section 3 where we develop the analytic apparatus for constructing globally optimal Wannier functions. After this are Sections 4 and 5 where we describe our numerical procedure. Our numerical procedure is illustrated in Section 6 which contains numerical experiments. We provide our conclusions and outline future work in Section 7.
@@ 2.1 Notation
We let $\ell^{2}(\mathbb{Z})$ denote the Hilbert space for complex sequences $\mathbf{x}=\left\{x_{i}\right\}_{i=-\infty}^{\infty}$ such that $\sum_{i=-\infty}^{\infty}\absolutevalue{x_{i}}^{2}<\infty$ with the standard inner product. We denote the right shift operator in $\ell^{2}(\mathbb{Z})$ by $R$, where
@@
For a vector $\mathbf{x}\in\ell^{2}(\mathbb{Z})$ or $\mathbb{C}^{N}$, we denote by $\mathbf{x}^{*}$ the conjugate transpose and by $\|\mathbf{x}\|$ the standard 2-norm, $\|\mathbf{x}\|=\sqrt{\mathbf{x}^{*}\mathbf{x}}$.
Given sets $A\subset\mathbb{C}$ and $B\subset\mathbb{C}$, we write $f\in C(A;B)$ if $f:A\to B$ is a continuous function. Similarly, we write $f\in C^{n}(A;B)$ if $f$ is a $n$-times continuously differentiable. For an $\mathbf{A}$ as an operator in $\ell^{2}(\mathbb{Z})$ or a matrix in $\mathbb{C}^{M\times M}$, we let $\mathbf{A}^{\dagger}$ denote its Moore–Penrose pseudoinverse.
@@ 2.2 Bloch functions
$\psi_{k}^{(j)}(x+y)=e^{\mathrm{i}ky}\psi_{k}^{(j)}(x),\quad y\in\Lambda.$ (10)
The value $k$ is called the quasimomentum, $\{\psi_{k}^{(j)}\}_{j=1}^{\infty}$ in (9) are called the Bloch functions, and $\{E_{k}^{(j)}\}_{j=1}^{\infty}$ are called the energy of the states represented by $\{\psi_{k}^{(j)}\}_{j=1}^{\infty}$.
@@
$\left[-\left(\frac{\mathrm{d}}{\mathrm{d}x}+\mathrm{i}k\right)^{2}+V\right]u_{k}^{(j)}=E_{k}^{(j)}u_{k}^{(j)},$ (13)
@@
$H(k)=-\left(\frac{\mathrm{d}}{\mathrm{d}x}+\mathrm{i}k\right)^{2}+V.$ (14)
@@ 2.3 Wannier functions
$W_{0}^{(j)}(x)=\frac{1}{\Omega}\int_{-\Omega/2}^{\Omega/2}\psi_{k}^{(j)}(x)\,{\rm d}k=\frac{1}{\Omega}\int_{-\Omega/2}^{\Omega/2}e^{\mathrm{i}xk}u_{k}^{(j)}(x)\,{\rm d}k,\quad j=1,2,\ldots.$ (22)
Wannier functions are highly non-unique since the Bloch functions in (9) with the normalization condition in (16) is only determined up to a complex number of unit modulus. More explicitly, for any $\psi_{k}^{(j)}$ satisfying (9), so will any $e^{-\mathrm{i}\phi(k)}\psi_{k}^{(j)}$ with a real $\phi(k)$ for $k\in I_{\mathrm{bz}}$ without affecting the normalization (16). In this paper, we refer to numbers of the form $e^{-\mathrm{i}\phi}$ as phase factors. For both physical and computational purposes, it is highly desirable to choose $\psi_{k}^{(j)}$ so that the Wannier functions are well-localized.
It was shown in [14], in the case of a symmetric potential, and in [9], more generally, that $\{\psi_{k}^{(j)}\}_{j=1}^{\infty}$ can be chosen such that the $\{W_{0}^{(j)}\}_{j=1}^{\infty}$ are exponentially localized. In other words, there exist constants $C>0$ and $D>0$ such that
@@
The variance defined by $\langle x^{2}\rangle-\langle x\rangle^{2}$ is a standard localization measure for Wannier functions [20].
@@ 2.4 Analyticity of eigenvalues and eigenvectors
Theorem 2.3 concerns only the eigenprojectors, not the eigenvectors; this is because for any an eigenvector $\mathbf{v}_{i}(z)$ of $T(z)$, one can multiply the vector by a $z$-dependent phase factor $e^{-\mathrm{i}\phi(z)}$ ($\phi(z)$ is real for real $z$) while the new vector $e^{-\mathrm{i}\phi(z)}\mathbf{v}_{i}(z)$ is still an eigenvector. The phase factor could change the smoothness of $\mathbf{v}_{i}(z)$ while the projector $P_{i}(z)=\mathbf{v}_{i}(z)\mathbf{v}^{*}_{i}(z)$ is unchanged for real $z$.
@@ 2.5 Derivatives of eigenvalues and eigenvectors
Next, we provide expressions for the derivatives of eigenvalues and eigenvectors. For simplicity, we only state the result for the Hilbert space $\ell^{2}(\mathbb{Z})$ and operators with a purely discrete spectrum, which will suffice for our purpose. These formulas can be found in Section 8.2.3 in [13]. The formula (28) is a consequence of the derivative of the projection $\frac{\mathrm{d}P}{\mathrm{d}z}$ substituted into (26).
@@
Suppose that $T(z)$ is a self-adjoint analytic family of operators in $\ell^{2}(\mathbb{Z})$ for $z\in D$ and the operators have a purely discrete spectrum. Suppose further that $z_{0}\in D$ is real and $\lambda(z_{0})$ is a nondegenerate eigenvalue of $T(z_{0})$ with the eigenvector $\mathbf{v}$, i.e. $T(z_{0})\mathbf{v}(z_{0})=\lambda(z_{0})\mathbf{v}(z_{0})$. The derivative of $\lambda$ at $z_{0}$ is given by
@@ 3.1 Analysis of Bloch functions
$u(x,k+g)=e^{-\mathrm{i}gx}u(x,k),\quad g\in\Lambda^{*},\,x\in I_{\mathrm{uc}}.$ (34)
@@
The following lemma provides a simple relation for the dependence of $k$ and $m$ in the Fourier coefficients in (35). It is a well-known relation in the context of periodic zone scheme (see (17)-(19)). It is a simple consequence of (34).
@@
$a_{m}(k)=\alpha(k+m\Omega),\quad k\in I_{\mathrm{bz}},\,m\in\mathbb{Z}.$ (39)
@@
Lemma 39 shows that we can obtain the function $\alpha$ defined in $\mathbb{R}$ by “unfolding” the Fourier coefficients $a_{m}(k)$ defined in $I_{\mathrm{bz}}$ for $m\in\mathbb{Z}$ such that $a_{m}(k)$ takes up the interval $[-\Omega/2+m\Omega,\Omega/2+m\Omega)$. We summarize this fact in the following lemma, together with Lemma 3.3 and the decay condition (38) in terms of $\alpha$.
@@
where $\alpha$ is defined in (39). The function $\alpha(k)$ can be chosen to be analytic in the intervals $[-\Omega/2+m\Omega,\Omega/2+m\Omega)$ for $m\in\mathbb{Z}$. Furthermore, |$\absolutevalue{\alpha(k)}$ | decays no slower than $|k|^{-2}$ as $|k|\to\infty$.
@@
Since $u(x,k)$ may not be periodically analytic in $k\in I_{\mathrm{bz}}$ (see Remark 3.2), the function $\alpha$ can be a discontinuous function in $\mathbb{R}$. Lemma 3.3 only ensures $\alpha$ is analytic in each interval $[-\Omega/2+m\Omega,\Omega/2+m\Omega)$ for $m\in\mathbb{Z}$. Moreover, the function $\alpha$ is formed from a very special choice of coefficients $a_{m}$ since the corresponding $u(x,k)$ is chosen to be analytic in $k$; this does not happen if one computes $u(x,k)$ at each point $k$ independently (see Remark 2.5) so that $u(x,k)$ is an arbitrarily irregular function in $k$.
@@ 3.2 Properties of Wannier functions
$W_{0}(x)=\frac{1}{\Omega}\int_{-\Omega/2}^{\Omega/2}\sum_{m=-\infty}^{\infty}\alpha(k+m\Omega)e^{\mathrm{i}(k+m\Omega)x}\mathrm{d}k=\frac{1}{2\pi}\int_{-\infty}^{\infty}\frac{2\pi}{\Omega}\alpha(\xi)e^{\mathrm{i}\xi x}\mathrm{d}\xi.$ (47)
@@
$\widehat{W}_{0}(\xi)=\int_{-\infty}^{\infty}W_{0}(x)e^{-\mathrm{i}\xi x}\,{\rm d}x=\frac{2\pi}{\Omega}\alpha(\xi).$ (49)
@@ 3.3 Perturbation analysis in the Fourier domain
where the vector with elements $\{\widehat{V}_{l}\}_{l=-\infty}^{\infty}$ is in $\ell^{2}(\mathbb{Z})$ for piecewise continuous $V$ in $I_{\mathrm{uc}}$. Moreover, we define a vector $\mathbf{y}$ containing the Fourier coefficients of $u$ in (35) and (45):
@@
It follows that $\mathbf{y}\in\ell^{2}(\mathbb{Z})$ for any $u$ that satisfies the assumptions of Lemma 3.5. Moreover, it follows from (37) that
@@
We observe that the operator $-\mathbf{\Theta}^{\dagger}(k)\mathbf{S}(k)$ in (61) has a spectrum that decays as $1/m$ asymptotically for any $k\in I_{\mathrm{bz}}$, so it is a compact operator on $\ell^{2}(\mathbb{Z})$. As a result, despite the fact that (61) is infinite-dimensional, (61) is very benign both mathematically and numerically.
@@ 3.4 Analysis of the discontinuities at zone boundaries
In this section, we show that the discontinuities of the Fourier transform $\alpha(k)$ (and its derivatives of all orders) at $k=\Omega/2+m\Omega$ for $m\in\mathbb{Z}$ are identical.
@@
As discussed in Remark 3.6, $a_{m}(\Omega/2)$ is not guaranteed to be equal to $a_{m+1}(-\Omega/2)$. Since $a_{m}(\Omega/2)$ and $a_{m+1}(-\Omega/2)$ may be viewed as limits of $\alpha(k)$ at the boundary points $k=\Omega/2+m\Omega$ for $m\in\mathbb{Z}$ from both sides, their differences characterize the jump of $\alpha(k)$ at these boundary points. The following lemma shows that $a_{m}(\Omega/2)$ and $a_{m+1}(-\Omega/2)$ only differ by a phase factor independent of $m$.
@@ 3.5 Gauge transformation
In this section, we introduce gauge transformation and show that a simple class of gauge choices fixes all discontinuities of the Fourier transform $\alpha$ discussed in Section 3.4, thus obtaining exponentially localized Wannier functions.
@@ 3.7 Realty of Wannier functions
While we have completed the construction of optimally localized Wannier functions, there is one remaining degree of freedom, the phase choice of the vector $\mathbf{y}_{0}$ in (63); we could replace $\mathbf{y}_{0}$ by $e^{-\mathrm{i}\phi_{0}}\mathbf{y}_{0}$ for any real $\phi_{0}$ and all above formulas would still hold, including the expressions in Corollary 3.14. However, the phase factor $e^{-\mathrm{i}\phi_{0}}$ does change the resulting $W_{0}$. In this section, we prove that the constructed Wannier function $W_{0}$ can always be chosen to be real if $\mathbf{y}_{0}$ in (63) satisfies a simple condition.
@@
We observe that the condition in (100) is applicable after solving (61) for determining $\mathbf{y}$ in $[-\Omega/2,\Omega/2]$; we can obtain a constant phase factor $e^{-\mathrm{i}\phi_{0}}$ to enforce (100) so the the resulting $W_{0}$ is real.
@@ 6 Numerical experiments
We implemented the procedure described in Section 4 in Fortran. We fix $V$, $a$, $M$, and $\ell$ and increase $K$ until 10 digits of accuracy is obtained. The timings reported are the CPU times required to obtain the Fourier transform of $W_{0}$ evaluated on a uniform grid (i.e., Steps 1–5 in Section 5). To obtain an estimate of the error we compute
@@ 7 Conclusion and generalization
Since Kato’s theorem (Theorem 2.3) guarantees analyticity even when eigenvalues become degenerate in one dimension, the approach in this paper has been easily extended to isolated multiband cases in one dimension. Similarly, globally optimal solutions can be computed analytically. For higher dimensions, topological phenomena show up in the construction of Wannier functions and new ideas beyond this paper are required. Nonetheless, the algorithms in one dimension are the basic building block for higher-dimensional cases. Such extensions to higher dimensions have been worked out and are now in preparation for publication.